OpenStudy (anonymous):

What is the remainder when is divided by 10?

OpenStudy (gorv):

what is divided by 10??

OpenStudy (perl):

any number divided by 10, the remainder is the units place , or ones place digit

OpenStudy (perl):

so for example 138 / 10 , the remainder is 8

OpenStudy (anonymous):

what about 4 to 31st power ?

OpenStudy (perl):



OpenStudy (perl):

you want the remainder when dividing by 10 ?

OpenStudy (anonymous):

yes

OpenStudy (akashdeepdeb):

\(4^{31}\) 4^1 = 4 4^2 = 16 4^3 = 64 4^4 = 256 4^5 = 1024 . . . All odd powers give units place => 4 All even powers give units place => 6 Hence, 31, being odd, will give remainder as 4 when divided by 10.

OpenStudy (perl):

basically theres a pattern when you take 4^n modulo 10 , you get alternating 4,6,4,6, ...

OpenStudy (perl):

for n>=1

OpenStudy (perl):

@AkashdeepDeb this is an inductive argument. not a 'proof' or deductive argument

OpenStudy (perl):

there could be a way to prove it using modular math consider 4^(2k) and 4^(2k+1)

ganeshie8 (ganeshie8):